Find the center of mass for each of the following regions.

1. The region bounded by y = x^3 ,x = -2 and the x-axis.


1
Expert's answer
2022-04-29T09:23:55-0400
"A=-\displaystyle\int_{-2}^{0}x^3dx=-[\dfrac{x^4}{4}]\begin{matrix}
 0 \\
 -2
\end{matrix}=4"

"\bar{x}=\dfrac{1}{4}(-\displaystyle\int_{-2}^{0}x(x^3)dx)=-\dfrac{1}{4}[\dfrac{x^5}{5}]\begin{matrix}
 0 \\
 -2
\end{matrix}=-\dfrac{8}{5}"

"\bar{y}=\dfrac{1}{4}(-\dfrac{1}{2}\displaystyle\int_{-2}^{0}(x^3)^2dx)=-\dfrac{1}{8}[\dfrac{x^7}{7}]\begin{matrix}
 0 \\
 -2
\end{matrix}=-\dfrac{16}{7}"

The centroid is "(-\dfrac{8}{5}, -\dfrac{16}{7})."